OpenStudy (anonymous):

Product: -42 , Sum: -1 [ :

OpenStudy (amistre64):

trying to find 2 numbers to match this?

OpenStudy (anonymous):

Yess [ :

OpenStudy (amistre64):

you want the easy answer or the long drawn out answer :)

OpenStudy (anonymous):

Easy lol

OpenStudy (amistre64):

lol ...... 7 times 6 = 42, one of thems gotta be a negative, id go with the 7

OpenStudy (anonymous):

Not sure if serious.

OpenStudy (amistre64):

ab = -42 a+b = -1 a = -1-b -42 = (-1-b)(b) -42 = -b -b^2 b^2 +b - 42 = 0 (b+7)(b-6) = 0 b = -7 or b = 6 this is actually both answer; -7(6) = 42 -7 + 6 = -1 thats the long drawn out answer ;)

OpenStudy (anonymous):

:o Why Thank You . [ ;